# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v2

@@ VII-C Communication Waveform KLD-based Optimisation
Fig. 7 illustrates the performance of the radar and communication subsystems using the optimised communication waveform. Fig. 7.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for power allocations of $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$. As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$. For example, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$, increasing to $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The linear improvement in $\mathrm{KLD_{r}}$ beyond $15$ dB indicates stable radar performance despite the suboptimal waveform. Fig. 7.b depicts the communication subsystem performance with $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$. $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$, rising to $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem thresholds with communication enhancements across different power allocations.
@@
Figure 7: The performance of both subsystems using optimised communication waveform. a) $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$; b) $\mathrm{KLD_{\mathrm{c}}}$ versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 8.a shows the ROC curves for the three targets at different $P_{\mathrm{R}}/N_{o}$ levels. At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.217,0.109,0.273\}$ for the non-optimal case. These results indicate that communication optimisation enhances detection capabilities at lower $P_{\mathrm{R}}/N_{o}$ compared to non-optimisation, as shown in Fig. 2. Fig. 8.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$. The BER decreases with increasing $P_{\mathrm{R}}/N_{o}$, reflecting improved communication performance. $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. In the non-optimal scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB (Fig. 1), while with optimisation, $P_{e}$ drops below $10^{-4}$ by 15 dB across all UEs. Although optimised waveforms provide a performance gain over the non-optimal method, the gain is less pronounced than with optimised radar waveforms due to radar interference.
Figure 8: a) The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. b) The BER versus $P_{\mathrm{R}}/{N}_{o}$.
Fig. 9.a illustrates the trade-off between both subsystems using the optimised communication waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values. As $P_{\mathrm{R}}/N_{o}$ increases from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems. This highlights the robustness of the optimisation technique in balancing radar and communication performance. In contrast, the non-optimal case in Fig. 3.b shows $\mathrm{KLD_{c}}$ saturating around 11 bits. Fig. 9.b presents the average CPU time for the communication waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$. The computational complexity stabilises after $15$ dB as the ISAC system transitions from an infeasible region, where minimum thresholds cannot be met for all targets and UEs, to a feasible region, with average CPU time fluctuating between $0.02$ and $0.04$ seconds.
Figure 9: a) The trade-off performance between the two subsystems $\mathrm{KLD_{\mathrm{c}}}$ versus $\mathrm{KLD_{\mathrm{r}}}$. b) The average CPU time (seconds) versus $P_{\mathrm{R}}/{N}_{o}$.
VIII Conclusion
@@
This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems. The two considered optimisation techniques, namely, KLD-based radar waveform optimisation and KLD-based communication precoding optimisation, demonstrated significant improvements over non-optimised scenarios. Radar waveform KLD-based optimisation yielded substantial enhancements in both target detection and communication performance, as evidenced by the ROC curves and reduced BER across all UEs. Communication KLD-based precoding optimisation, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs. Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimisation demonstrating stable computational efficiency. Moreover, trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances the overall ISAC system performance. These findings provide valuable insights for designing and optimising shared deployment ISAC systems, offering a promising approach to balance sensing and communication requirements in future wireless networks. The proposed framework provides the means for a holistic ISAC design, paving the way for more efficient and versatile wireless communications and sensing technologies.
@@ Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimisation problem.
A-A Gradients for Radar-Waveform Optimisation
The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (39), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (36). Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$. Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, we obtain,
$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}=\mathbf{R}_{2,t}^{-1}$ (58)
For the trace term, applying $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ yields,
$\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}(\sigma_{n}^{2}\mathbf{I}_{N}))}{\partial\mathbf{R}_{2,t}}=-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}$ (59)
Applying the chain rule to (36), and substituting (58) and (59), we get,
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}}$ (60)
@@ A-A Gradients for Radar-Waveform Optimisation
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{d_{\mathrm{r},t}^{\zeta}}{L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ -\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t})$ (61)
@@
$[\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})]_{i,l,t}=\frac{2\>d_{\mathrm{r},t}^{\zeta}}{TL\ln{2}}(\mathbf{H}_{t}^{H}(\mathbf{R}_{2,t})^{-1}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t}\\ -\sigma_{n}^{2}(\mathbf{R}_{2,t})^{-2}\mathbf{H}_{t}\mathbf{W}_{\mathrm{r},t})_{i}$ (62)
@@
$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}$ (64)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}$ (65)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}$ (66)
@@
$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}$ (67)
Equations (62) and (67) provide the mathematical foundation for the gradients used in our optimization algorithm.

### v4

@@ IX-C Communication Waveform KLD-based Optimisation
Figure 3: The ISAC system performance at $P_{\mathrm{r}}=0.5$ using communication waveform optimisation: K-COP. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
Fig. 3 illustrates the performance of both subsystems using the KLD-based communication waveform optimisation (K-COP) technique. In Fig. 3.a, the achievable $\mathrm{KLD_{\mathrm{r}}}$ (right axis) and detection probability $P_{D}$ (left axis) versus $P_{\mathrm{R}}/N_{o}$ are shown for each target. For K-COP, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{\mathrm{r}}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{\mathrm{r}}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{10.12,8.13,10.91\}$ bits, increasing to $\{23.84,21.1,24.85\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The linear improvement in $\mathrm{KLD_{\mathrm{r}}}$ beyond $15$ dB for both techniques indicates stable radar performance despite the suboptimal waveform. At $P_{\mathrm{R}}/N_{o}=18$ dB, the $P_{D}$ values for K-COP are $\{72.2,60.7,75.4\}\%$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ compared to $\{61,48.3,64.2\}\%$ for the non-optimised ZF shown in Fig. 1. K-COP demonstrates improved detection capabilities over non-optimisation but remains below the radar waveform optimisation performance achieved using K-ROP in Fig. 2, which is $\{98,89.2,99.9\}\%$.
Fig. 3.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{\mathrm{c}}}$ (right axis) and BER (left axis) versus $P_{\mathrm{R}}/N_{o}$. For K-COP, $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{\mathrm{c}}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. For instance, at $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{36.94,19.55,66.52\}$ bits, rising to $\{87.06,71.15,95.94\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. In the non-optimised ZF scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after $15$ dB (Fig. 1), while with K-COP optimisation, $P_{e}$ drops below $10^{-5}$ by $14$ dB across all UEs. K-COP optimisation technique provides significant performance gains over the non-optimised ZF scenario, though these gains are less pronounced than those achieved through radar waveform optimisation in Sec. IX-B due to radar interference. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem performance with communication enhancements.
IX-D Integrated Waveform KLD-based Optimisation for ISAC
The integrated waveform optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, $\rho=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
Figure 4: The ISAC system performance using the integrated waveform optimisation. a) Radar subsystem performance: $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ vs $P_{\mathrm{R}}/N_{o}$. b) Communication subsystem performance: BER and $\mathrm{KLD}_{\mathrm{c}}$ vs $P_{\mathrm{R}}/{N}_{o}$.
Fig. 4 illustrates the ISAC system performance under the KLD-based integrated waveform optimisation (K-IOP). In Fig. 4.a, the radar subsystem performance is shown in terms of detection probability $P_{D}$ (left axis) and $\mathrm{KLD}_{\mathrm{r}}$ (right axis) for each target $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$. As $\mathit{P}_{\mathrm{R}}/N_{o}$ increases from 0 to 30 dB, both $P_{D}$ and $\mathrm{KLD}_{\mathrm{r}}$ improve for all targets, confirming the effectiveness of K-IOP. For instance, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, the achievable $\mathrm{KLD}_{\mathrm{r}}$ values for $\{\mathrm{T}_{1},\mathrm{T}_{2},\mathrm{T}_{3}\}$ are approximately $\{38.87,32.77,41.85\}$ bits, with corresponding detection probabilities $\{0.94,0.89,0.97\}$. Performance continues to rise at higher $\mathit{P}_{\mathrm{R}}/N_{o}$, achieving near-unity detection across all targets once $\mathit{P}_{\mathrm{R}}/N_{o}$ exceeds 22 dB.
In Fig. 4.b, the communication subsystem performance is depicted via the BER (left axis) and $\mathrm{KLD}_{\mathrm{c}}$ (right axis) for each UE $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$. Similar to the radar trends, higher $\mathit{P}_{\mathrm{R}}/N_{o}$ values yield a marked improvement in both $\mathrm{KLD}_{\mathrm{c}}$ and BER. For example, at $\mathit{P}_{\mathrm{R}}/N_{o}=10$ dB, $\mathrm{KLD}_{\mathrm{c}}$ values of $\{77.19,28.13,260.53\}$ bits are observed for $\{\mathrm{UE}_{1},\mathrm{UE}_{2},\mathrm{UE}_{3}\}$, with corresponding BERs of approximately $\{5.1\times 10^{-9},1.7\times 10^{-4},\lesssim 10^{-10}\}$. Beyond 14-16 dB, the BER for all UEs drops rapidly– often reaching effectively error-free communication– and $\mathrm{KLD}_{\mathrm{c}}$ grows significantly as the interference from the radar subsystem is jointly minimised. Comparing these K-IOP results with the non-optimised ZF baseline (cf. Fig.1) or the separate (radar-only or communication-only) waveform approaches in Sec. IX-BIX-C, it is evident that the integrated design offers substantially improved performance simultaneously for radar detection and communication reliability. This highlights the advantages of a truly integrated ISAC paradigm, where waveforms are optimally tailored to balance both subsystems’ requirements.
Figure 5: Average MSE of the target response matrix versus $P_{\mathrm{R}}/N_{o}$ for all optimisation techniques.
Fig. 5 presents the average MSE of the estimated target response matrix $\mathbf{\hat{H}}_{t}$ averaged for all targets versus $P_{\mathrm{R}}/N_{o}$ for K-ROP, K-COP, and K-IOP optimisation techniques. At low $P_{\mathrm{R}}/N_{o}$ values (00-$10$ dB), K-IOP immediately achieves lower MSE than both K-ROP and K-COP. As $P_{\mathrm{R}}/N_{o}$ increases beyond 10 dB, the techniques begin to show more distinct performance. K-ROP demonstrates faster MSE decay, particularly in the $15$-$20$ dB range, while K-COP exhibits a steadier improvement but converges slightly more slowly. The K-IOP continues to outperform both separate waveform approaches at all values of $P_{\mathrm{R}}/N_{o}$, reaching near-zero MSE at lower $P_{\mathrm{R}}/N_{o}$ values. Eventually, all three techniques attain negligible MSE at sufficiently high $P_{\mathrm{R}}/N_{o}$ levels. Notably, this MSE performance strongly correlates with the detection capabilities in Figs. 3.a, 2.a, and 4, where more reliable detections yield more accurate target-response estimates under $\mathcal{H}_{1}$. Conversely, missed detections or false alarms degrade the estimation by introducing noise-based or zero-valued entries.
Figure 6: Average CPU time (seconds) versus antenna array size $N$ for both optimisation techniques.
Fig. 6 presents the average CPU time versus antenna array size $N$, comparing all optimisation techniques. The K-ROP consistently demonstrates the lowest computational cost, with CPU times increasing from $0.001$ s at $N=10$ to $7.98$s at $N=100$. Meanwhile, K-COP exhibits moderate requirements, ranging from $0.04$ s to $14.07$s across the same $N$. By contrast, the K-IOP incurs the highest computational load, growing from $2.44$ s at $N=10$ to $37.69$s at $N=100$. For instance, at $N=50$, K-ROP requires only $2.80$s, whereas K-COP needs $3.90$s, and the K-IOP demands $19.96$s. These results indicate that while the radar-only and communication-only algorithms remain relatively efficient as $N$ increases, the K-IOP entails modestly higher computational complexity, while offering much higher performance.
X Conclusion
@@
This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems. The three considered optimisation techniques, namely, K-ROP, K-COP, and K-IOP for ISAC approach, each demonstrated significant improvements over the non-optimised scenario. K-ROP yielded substantial enhancements in both target detection and communication performance while maintaining efficient computational requirements, as evidenced by the numerical results. K-COP, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower $P_{\mathrm{R}}/N_{o}$ levels. Both techniques exhibited robust performance across varying SNR values. In addition, the K-IOP effectively unifies these advantages, consistently outperforming the two separate approaches in terms of detection probability, BER, and estimation accuracy. By jointly designing radar and communication waveforms, the K-IOP approach leverages the synergy between sensing and data transmission, thereby achieving higher overall KLD and a more balanced performance for future ISAC networks. Moreover, the numerical results revealed that exploiting the communication signal as part of the radar subsystem significantly boosts ISAC system capabilities, underscoring the importance of holistic design in shared-antenna deployments. These findings provide valuable insights for designing and optimising ISAC systems, offering a promising pathway to balance sensing and communication requirements in next-generation wireless networks. The proposed framework offers the flexibility to adapt system resources for diverse operating regimes, paving the way for more efficient and versatile wireless communications and sensing technologies.
@@ Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimisation problem. The objective function $f(\mathbf{\dot{W}}_{\mathrm{r}})$ is given in (31), where the $\mathrm{KLD}_{\mathrm{r},t}$ is defined in (27). Therefore, to derive $\nabla f(\mathbf{\dot{W}}_{\mathrm{r}})$, we first compute $\partial\mathrm{KLD}_{\mathrm{r},t}/\partial\mathbf{W}_{\mathrm{r},t}$. Using the matrix identity $\frac{\partial\ln(\det(\mathbf{X}))}{\partial\mathbf{X}}=(\mathbf{X}^{-1})^{T}$, and derivative $\frac{\partial\mathrm{Tr}(\mathbf{X}^{-1}\mathbf{A})}{\partial\mathbf{X}}=-(\mathbf{X}^{-1})^{T}\mathbf{A}^{T}(\mathbf{X}^{-1})^{T}$ we obtain,
$\frac{\partial\ln(\det(\mathbf{R}_{2,t}))}{\partial\mathbf{R}_{2,t}}\!=\!\mathbf{R}_{2,t}^{-1},\quad\frac{\partial\mathrm{Tr}(\mathbf{R}_{2,t}^{-1}\sigma_{n}^{2})}{\partial\mathbf{R}_{2,t}}\!\!=\!\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}.$ (68)
Applying the chain rule to (27), and substituting (68) we get,
$\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\mathrm{r},t}}=\frac{1}{\ln{2}}(\mathbf{R}_{2,t}^{-1}-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2})\cdot\frac{\partial\mathbf{R}_{2,t}}{\partial\mathbf{W}_{\mathrm{r},t}},$ (69)
@@
$\!\!\!\!\!\frac{\partial\mathrm{KLD}_{\mathrm{r},t}}{\partial\mathbf{W}_{\!\mathrm{r},t}}\!\!=\!\!\frac{\!d_{\mathrm{r},t}^{\zeta}}{\!L\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}-\sigma_{n}^{2}\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!).\!\!$ (70)
@@
$\!\!\!\!\!\![\nabla\!f(\!\mathbf{\dot{W}}\!_{\mathrm{r}}\!)]_{i,l,t}\!\!=\!\!\frac{2d_{\mathrm{r},t}^{\zeta}}{TL\!\ln{2}}(\mathbf{H}_{t}^{H}\mathbf{R}_{2,t}^{-1}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t}\!-\sigma_{n}^{2}\mathbf{R}_{2,t}^{-2}\mathbf{H}_{t}\!\mathbf{W}_{\!\mathrm{r},t})_{i}.\!\!\!$ (71)
@@
$\frac{\partial\sigma_{\eta}^{2}}{\partial\mathbf{W}_{\mathrm{r},l}}=\frac{2\>d_{\mathrm{c},k}^{\zeta}\sigma_{h}^{2}}{L}\mathbf{W}_{\mathrm{r},l}.$ (73)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\sigma_{\eta}^{2}}=-\frac{\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)}{M(M-1)\sigma_{\eta}^{2}\ln 2}.$ (74)
@@
$\frac{\partial\mathrm{KLD}_{\mathrm{c},k}}{\partial\mathbf{W}_{\mathrm{r},l}}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}\mathbf{W}_{\mathrm{r},l}.$ (75)
@@
$[\nabla\mathrm{KLD}_{\mathrm{c},k}]_{i,l,t}=-\frac{2\lambda\>d_{\mathrm{c},k}^{\zeta}p_{k}(N-K)\sigma_{h}^{2}}{M(M-1)L\sigma_{\eta}^{2}\ln 2}[\mathbf{\dot{W}}_{\mathrm{r}}]_{i,l,t}.$ (76)
Equations (71) and (76) provide the mathematical foundation for the gradients used in our optimisation algorithm.